# Allocating Positional Goods: A Mechanism Design Approachthanks: I thank Krishna Dasaratha, Alex Gershkov, Bart Lipman, Chiara Margaria, and Juan Ortner for their helpful comments. The paper subsumes results in my note “A Mechanism Design Approach to ‘Gainers and Losers in Priority Services’.”

### v2

@@ 4.3 Status Game As a Suffering
In this subsection, I consider the case where $v^{\prime}(\theta)\leq-1$, so the status game is a “suffering” in the sense that participants with higher types suffer more and receive lower payoffs. In the queuing example (Example 2.1), $v^{\prime}(\theta)\leq-1$ holds if higher types derive lower utility from the service after accounting for waiting costs—for example, if the service has the same value $v_{0}$ to everyone: $u(p,s,\theta)=v_{0}-\theta t-p$, where $t=1-s$ is the waiting time (and thus $v(\theta)=v_{0}-\theta$). Because higher types face higher costs or disutility, agents have incentives to overreport their types to get lower waiting time.
@@
    Assume $L(\theta)=\theta+F(\theta)/f(\theta)$ is increasing. Then,
@@
      If $-v^{\prime}(\theta)/v(\theta)\leq f(\theta)/F(\theta)$ for all $\theta\in[0,\bar{\theta}]$, then no exclusion is optimal (i.e., $\theta_{0}^{*}=\bar{\theta}$).
@@
    The seller can obtain at least half the maximum revenue by selling one level of positional good.
@@
Parts (i)–(iii) are qualitatively similar to those in the main specification, except that the revenue-maximizing mechanism excludes high types instead of low types. For part (iv), when the seller assign negative status, she has to assign it to low types first because $s$ is increasing (by incentive compatibility). Hence, because $U^{\prime}(\theta)\leq 0$, negative status reduces the payoffs of all participants, and therefore cannot increase the seller’s revenue.
@@ 5 Conclusion
In this paper, I study the allocation of positional goods using a mechanism approach. Many goods are positional in the sense that allocating an item to one agent has externalities on others: moving up the position of one consumer would push down that of another, and the value of the good depreciates as more people have it. The model has dual implications in the monopolist provision of positional goods and the design of status categories within organizations. The results also shed light on the trade-offs involved in educational disarmament.
@@
In terms of the monopolist problem, the seller can often extract more revenue by offering many levels of positional goods through their status values. At the extreme, an all-pay auction with reserve prices, where the more buyers pay, the higher positions they receive, can maximize the seller’s revenue. On the other hand, offering more levels of positional goods is often detrimental to consumers. Thus, it is usually beneficial for consumers to restrict the number of position levels the seller can offer, and the consumer-optimal regulation without using subsidies is to restrict the seller to only one levels of positional goods and serving everyone. Nevertheless, in cases where a few buyers have very high valuations (e.g., Pareto distribution), both the seller’s revenue and the consumer surplus are higher when more status levels are offered—their objectives are aligned.
@@
In terms of the organizational design problem, the results suggest that the management practice of stack ranking and rat races in education maximizes agents’ effort, consistent with the common wisdom, if the ability distribution satisfies Myerson’s regularity. The exclusion of low types can increase revenue or effort. On the contrary, from the agent’s perspective, less intense status competition is beneficial if the ability distribution satisfies IFR (or has a thin tail). This explains the call for abolishing stack ranking and educational disarmament, such as lotteries for school admissions. Nevertheless, when a few agents have very high abilities (e.g., Pareto distribution), rat races can be better for agent welfare. Thus, the objectives to maximize agent effort and welfare can be aligned, which challenges the common wisdom that rat races are always detrimental to agents. In either case, exclusion always harms agent welfare, even though it can increase the status of participants. Moreover, when learning effort is productive, education arms races can increase social welfare.
@@
It remains to be explored how competition among sellers affects the results. This can be interpreted as offering a different track of education that leads to different status levels (e.g., vocational versus academic education) or having another product that has a separate queue. In either case, consumers of one set of positional goods derive utility independently of those of the other, either because it is psychologically incomparable or physically separate. If the other seller offers the same menu as the original seller, an immediate result is that each seller will attract half of the buyers, and that competition can increase consumer surplus.242424 Di Stasio et al. (2016) find that education is less positional in countries with more developed vocational education systems (e.g., Germany). However, when sellers compete in mechanisms strategically, the problem remains an open question.
@@ Appendix A Optimal Mechanisms without Exclusion
In this section, I assume exclusion is impossible and that the lowest-level positional good is free. For example, in Moldovanu, Sela, and Shi (2007), agents cannot be excluded and will at least get the lowest status in the organization. In the benchmark model of Gershkov and Winter (2023), buyers can use the regular (non-priority) service for free. Because exclusion is not allowed, the free low-level position, rather than opting out, becomes the relevant outside option. As noted in Remark 2, the feasibility condition becomes $s\in\operatorname{\text{MPS}}(F)$. I will repeat the analysis in Section 3 and show that the results are qualitatively similar.
@@ A.1 Revenue Maximization
The revenue is given by
@@
By (IR), it is optimal to set $U(0)=v(0)$. Denote $J(\theta)=\theta-\frac{1-F(\theta)}{f(\theta)}$. The revenue maximization problem is equivalent to
@@
Offering more positional good levels (strictly) increases the seller’s revenue if and only if $J(\theta)$ is (strictly) increasing. Thus, full separation maximizes revenue if and only if $J(\theta)$ is increasing.
@@
GW’s Proposition 8 provides a sufficient condition for the seller’s revenue to be strictly increasing in the number of priority classes: $F$ satisfies the IFR property (i.e., $\frac{1-F(\theta)}{f(\theta)}$ is decreasing).
@@
Effort maximization in MSS’s model is the same as revenue maximization because they assume linear effort costs (and no exclusion). They provide a sufficient condition for full separation in Theorem 4: $F$ satisfies the IFR property.
@@
{proofsketch}
Because $J^{\prime}(\bar{\theta})=2>0$, $J(\theta)$ is increasing in a neighborhood of $\bar{\theta}$, so by Proposition 2 in KMS, types in this neighborhood are separated.
@@
The seller can obtain at least half the maximum revenue by offering one positional good level in addition to a free low-level position.
@@
By offering one positional good level in addition to a free low-level position, the seller’s maximum revenue is
@@
Recall that $\tilde{J}(\theta)=\int_{0}^{\theta}J(x)\mathop{}\!\mathrm{d}F(x)=-(1-F(\theta))\theta$. Figure 4 plots the revenue curve $R(\tau)\equiv-\tilde{J}(F^{-1}(\theta))$ in the quantile space where $\tau=F(\theta)$. Assume $v(\theta)=0$ for simplicity. In Figure 4, (i) the area under the revenue curve (blue area in the left panel) represents the revenue from the revenue-maximizing mechanism—i.e., $s^{*}(\theta)=F(\theta)$, (ii) the red area in the right panel represents the revenue from having two levels of positional goods—a higher level offered at price $F(\theta^{*}_{0})$ in addition to a lower level offered for free, and (iii) the entire box of either panel (which has an area of $R(\tau^{*}_{0})$) represents the maximum revenue from selling nonpositional goods in the auxiliary screening problem. It is straightforward that the red area (ii) is larger than half of the entire box and thus strictly larger than half of the blue area (i).
@@
If $F(\theta)=1-\exp(-\lambda\theta)$ where $\lambda>0$, offering one positional good level (in addition to a free low level) can obtain $73.6\%$ of the maximum revenue.
@@
If $F(\theta)=\theta$ on $[0,1]$, offering one positional good level (in addition to a free low level) can obtain $75\%$ of the maximum revenue.
@@
If $F(\theta)=\theta^{\beta}$ on $[0,1]$ where $\beta>0$, offering one positional good level (in addition to a free low level) can obtain
@@ A.2 Consumer Surplus
If both $\theta-\frac{1-F(\theta)}{f(\theta)}$ and $\frac{1-F(\theta)}{f(\theta)}$ are increasing (e.g., exponential and Pareto distributions), a finer partition (in terms of majorization) can increase both the seller’s revenue and consumer surplus. Thus, full separation maximizes both the revenue and consumer surplus.
@@
$\int_{0}^{\theta}\left(\frac{1-F(x)}{f(x)}-\operatorname{\mathbf{E}}[\theta]\right)\mathop{}\!\mathrm{d}F(x)\geq 0\mbox{ for all $\theta\in[0,\bar{\theta}]$.}$
@@
Denote $I(\theta)=\int_{0}^{\theta}\frac{1-F(x)}{f(x)}\mathop{}\!\mathrm{d}F(x)$. Then, the condition in the proposition is equivalent to $I(\theta)\geq I(\bar{\theta})F(\theta)=\operatorname{\mathbf{E}}[\theta]F(\theta)$ (graphically, $I(\theta)$ lies above the line connecting $I(0)=0$ and $I(\bar{\theta})=\operatorname{\mathbf{E}}[\theta]$ in the quantile space). Therefore, $I(\bar{\theta})F(\theta)$, which corresponds to total pooling, is the convex hull of $I$ (i.e., the largest convex function that lies below $I$).252525 A piecewise affine function consisting of other points on $I(\theta)$, which corresponds to offering more positional good levels, lies above $I(\bar{\theta})F(\theta)$ and leads to lower consumer surplus. By Proposition 2 in KMS, this condition is necessary and sufficient for total pooling to be welfare-maximizing. ∎
@@
$\max_{s\in\operatorname{\text{MPS}}(F)}\int_{0}^{\theta}\frac{s(x)}{f(x)}\mathop{}\!\mathrm{d}F(x)+v(\theta)+U(0)-v(0).$ (A.9)
Because $p(0)\geq 0$, it is optimal to set $p(0)=0$ and $U(0)=v(0)$. The following proposition provides a necessary and sufficient condition for total pooling to maximize $U(\theta)$ for all $\theta\in[0,\bar{\theta}]$.
@@
Note that $\int_{0}^{\theta}1/f(x)\mathop{}\!\mathrm{d}F(x)=\theta$, and the condition $F(\theta)\leq\theta/\bar{\theta}$ is equivalent to $\int_{0}^{\theta}(1/f(x)-\bar{\theta})\mathop{}\!\mathrm{d}F(x)\geq 0$ for all $\theta\in[0,\bar{\theta}]$. Similar to the proof of Proposition A.5, one can show that this condition is necessary and sufficient for $s(\theta)=1/2$ to maximize $U(\theta)$ for all $\theta\in[0,\bar{\theta}]$.
@@ Appendix B Proofs
B.1 Proof of Section 3
@@
The following lemma follows from applying Theorem 4 (Fan-Lorentz) in KMS to seller’s revenue given $\theta_{0}$, that is,
@@
Note that under the specification $\tilde{S}({q},G_{\chi})=\phi(S({q},G_{\chi}))$, where $\phi$ is an increasing and convex function, $J(\theta)s(\theta)$ in equation (B.1) is replaced by $J(\theta)\phi(s(\theta))$. Lemma B.1 still follows from Fan-Lorentz Theorem because $J(\theta)\phi(s)$ is convex in $s$ and supermodular in $(s,\theta)$ when $\phi$ is increasing and convex.
On the other hand, if $J(\theta)$ is decreasing on $\Theta_{p}\subseteq[0,\bar{\theta}]$, Fan-Lorentz Theorem implies that the revenue-maximizing allocation pools all types $\theta\in\Theta_{p}$ at the same level. ∎
@@
Denote $J_{v}(\theta)=v(\theta)-\frac{1-F(\theta)}{f(\theta)}v^{\prime}(\theta)$. Under IFR, $J_{v}(\theta)$ is increasing because $v(\theta),v^{\prime}(\theta)\geq 0$ and $v^{\prime\prime}(\theta)\leq 0$. Denote
@@
Denote $\theta_{1}=J^{-1}(0)$. We have
@@
Denoted by $p$ the price of the positional good. The cutoff type $\theta(p)$ indifferent between buying and not buying is given by
@@
Because $v^{\prime}(\theta)\geq 0$, all types above $\theta(p)$ will buy the good. Thus, by selling one level of positional good, the seller’s maximum revenue is
@@
Consider the auxiliary screening problem of selling an indivisible item of value to one buyer, in which a standard result implies a posted-price mechanism is optimal (see Börgers, 2015, Proposition 2.5). Formally, denote by $q$ the probability the item is allocated to $\theta$, and let =$\mathcal{M}=\{q\colon[0,\bar{\theta}]\to[0,1]\mid q\text{ increasing}\}$. The agent’s payoff is $\theta q-p$. Then, the standard result implies that $\max_{q\in\mathcal{M}}\int_{0}^{\bar{\theta}}J(\theta)q(\theta)\mathop{}\!\mathrm{d}F(\theta)=\max_{p}p(1-F(p))$, and any maximizer $q^{*}$ must be an extreme point of $\mathcal{M}$ (i.e., a posted-price mechanism).
@@
(i) The following lemma follows from applying Theorem 4 (Fan-Lorentz) in KMS.
@@
If $s(\theta)=s_{0}$ is constant on $[\theta_{0},\theta_{1}]$ for some $\theta_{1}\in(\theta_{0},\bar{\theta}]$ (and $s(\theta)>s_{0}$ for $\theta>\theta_{1}$), then the consumer surplus is
@@
[…]
Denote
$z(\theta_{0})\equiv-\frac{W^{\prime}(\theta_{0})}{1-F(\theta_{0})}=v^{\prime}(\theta_{0})+F(\theta_{1})/2+F(\theta_{0})/2-\frac{f(\theta_{0})}{1-F(\theta_{0})}\int_{\theta_{0}}^{\theta_{1}}\frac{1-F(\theta)}{2}\mathop{}\!\mathrm{d}\theta.$
Note that $z(\theta_{1})-F(\theta_{1})-v^{\prime}(\theta_{1})=0$, and under IFR,
$z^{\prime}(\theta_{0})-f(\theta_{0})-v^{\prime\prime}(\theta_{0})=-\left[\frac{f(\theta)}{1-F(\theta)}\right]^{\prime}\int_{\theta_{0}}^{\theta_{1}}\frac{1-F(\theta)}{2}\mathop{}\!\mathrm{d}\theta\leq 0.$
Therefore, $z(\theta_{0})\geq F(\theta_{0})+v^{\prime}(\theta_{0})\geq 0$ and thus $W^{\prime}(\theta_{0})=-z(\theta_{0})(1-F(\theta_{0}))\leq 0$ for all $\theta_{0}\leq\theta_{1}$. Hence, decreasing $\theta_{0}$ increases the consumer surplus $W(\theta_{0})$. ∎
@@
[…]
It is straightforward that $W^{\prime}(\bar{\theta})=0$ and $W^{\prime\prime}(\bar{\theta})=(v^{\prime}(\theta_{0})+1)f(\bar{\theta})\geq 0$. Denote
$z(\theta_{0})\equiv-\frac{W^{\prime}(\theta_{0})}{1-F(\theta_{0})}=v^{\prime}(\theta_{0})+1/2+F(\theta_{0})/2-\frac{f(\theta_{0})}{1-F(\theta_{0})}\int_{\theta_{0}}^{\bar{\theta}}\frac{1-F(\theta)}{2}\mathop{}\!\mathrm{d}\theta,$ (B.8)
Note that $z(\bar{\theta})-F(\bar{\theta})-v^{\prime}(\bar{\theta})=0$, and under IFR,
$z^{\prime}(\theta_{0})-f(\theta_{0})-v^{\prime\prime}(\theta_{0})=-\left[\frac{f(\theta)}{1-F(\theta)}\right]^{\prime}\int_{\theta_{0}}^{\bar{\theta}}\frac{1-F(\theta)}{2}\mathop{}\!\mathrm{d}\theta\leq 0.$
Therefore, $z(\theta_{0})\geq F(\theta_{0})+v^{\prime}(\theta_{0})\geq 0$ and thus $W^{\prime}(\theta_{0})=-z(\theta_{0})(1-F(\theta_{0}))\leq 0$ for all $\theta_{0}>0$. Hence, no exclusion is optimal (i.e., $\theta^{*}_{0}=0$).

### v3

@@ 4.5 Status Game As a Suffering
Now I consider a case where $v^{\prime}(\theta)\leq-1$, so that the status game is a “suffering” in the sense that higher types suffer more and receive lower payoffs when they participate. In the queuing example (Example 1), if the good they are waiting for has homogenous values across agents, higher (cost) types will derive lower utility from it after accounting for waiting costs.  [^22] In Example 1, if the good has the same value $v_{0}$ to everyone, then the payoff $u(p,s,\theta)=v_{0}-\theta t-p=(v_{0}-\theta)+\theta s-p$ (because t=1st=1-s is the waiting time) satisfies $v^{\prime}(\theta)=-1$. In general, because higher types face higher costs or disutility, agents have incentives to overreport their types to get lower waiting time.
@@
    Assume $L(\theta)=\theta+\frac{F(\theta)}{f(\theta)}$ is increasing. Then,
@@
      If $\frac{v^{\prime}(\theta)}{v(\theta)}\leq\frac{f(\theta)}{F(\theta)}$ for all $\theta\in[0,\bar{\theta}]$, then no exclusion is optimal (i.e., $\theta_{0}^{*}=\bar{\theta}$).
@@
    The seller can obtain at least half the maximum revenue by selling a single good.
@@
Parts (i)–(iii) are qualitatively similar to those in the main specification, with the integrand $J(\theta)$ replaced by $L(\theta)=\theta+\frac{F(\theta)}{f(\theta)}$, and that the revenue-maximizing mechanism excludes high types instead of low types (because $U^{\prime}(\theta)\leq 0$). For part (iv), when the seller assigns a negative status, she must assign it to low types first because $s$ is increasing. However, since agents now have incentives to overreport (because $U^{\prime}(\theta)\leq 0$), assigning negative status to low types does not mitigate agents’ incentives to overreport and therefore cannot increase the seller’s revenue.
@@ 5 Conclusion
In this paper, I study the allocation of positional goods using a mechanism design approach. Many goods are positional because consumers care about their relative positions in consumption. Thus, allocating an item to one agent has externalities on others: moving up the position of one consumer would push down that of another, and the value of the good depreciates as more people have it. Because of these externalities, positional goods remain scarce and competitive despite economic growth.
@@
My model has implications for both the monopolist provision of positional goods and the design of status categories within organizations. In terms of the monopolist problem, the seller can often extract more revenue by offering many levels of positional goods through their status values. At the extreme, an all-pay auction with a reserve price, where the more a buyer pays, the higher position they receive, can maximize the seller’s revenue. By contrast, offering more levels of positional goods is often detrimental to consumers. Thus, it is usually beneficial for consumers to restrict the number of positional good levels the seller can offer, and the consumer-optimal regulation without using subsidies is to restrict the seller to offering a single good and serving everyone. Nevertheless, in cases where a few buyers have very high valuations (e.g., Pareto distribution), both the seller’s revenue and consumer surplus are higher when more status levels are offered—their objectives are aligned.
@@
The results also shed light on the trade-offs involved in educational disarmament. The results suggest that rat races in education maximize agents’ effort, consistent with the common wisdom, if the ability distribution satisfies Myerson’s regularity. By contrast, from the agent’s perspective, reduced status competition is beneficial if the ability distribution satisfies IFR (i.e., has a thin tail). This explains calls for educational disarmament, such as lotteries for school admissions. Nevertheless, when a few agents have very high abilities (e.g., Pareto distribution), educational arms races can be better for agents’ welfare. Intuitively, high-ability (low-ability) agents benefit from more (less) competition, so meritocracy raises (reduces) agents’ welfare on average when there are relatively more high-ability (low-ability) agents. Thus, the objectives of maximizing agents’ effort and welfare can be aligned, which challenges the common wisdom that rat races are always detrimental to agents. In either case, exclusion always harms agents’ welfare, even if it raises the status of participants. Similarly, in organizations, the results suggest that the management practice of stack ranking maximizes agents’ effort under Myerson’s regularity but harms agents’ welfare under IFR. Moreover, when learning effort is productive, education arms races can increase social welfare.
@@
It remains to be explored how competition among sellers affects the results. This can be interpreted as offering a different track of education that leads to different status levels (e.g., vocational versus academic education) or having another product that has a separate queue. In either case, consumers of one set of positional goods derive utility independently of those of the other, either because they are psychologically incomparable or physically separate. If the other seller offers the same menu, each seller will attract half of the buyers, and that competition can increase consumer surplus. [^23] Di Stasio, Bol and Van de Werfhorst (2016) find that education is less positional in countries with more developed vocational education systems (e.g., Germany). However, when sellers compete in mechanisms strategically, the problem remains open.
@@ Appendix A Optimal Mechanisms without Exclusion
In this section, I assume exclusion is impossible and that the lowest-level positional good is free. [^24] This section subsumes results in my note “A Mechanism Design Approach to ‘Gainers and Losers in Priority Services’.” For example, in Benny Moldovanu, Aner Sela and Xianwen Shi (2007), agents cannot be excluded and can get at least the lowest status in the organization. In the benchmark model of Gershkov and Winter (2023), buyers can use the regular (non-priority) service for free. Because exclusion is not allowed, the free low-level position, rather than opting out, becomes the relevant outside option. As noted in Remark 2, the feasibility condition becomes $s\in\operatorname{\text{MPS}}(F)$. I will repeat the analysis in Section 3 and show that the results are qualitatively similar.
@@ A.1 Revenue Maximization
By standard arguments, the expected revenue is given by
@@
By (IR), it is optimal to set $U(0)=v(0)$. The revenue maximization problem is equivalent to
@@
Offering more positional good levels always (strictly) increases the seller’s revenue if and only if $J(\theta)$ is (strictly) increasing. Thus, full separation maximizes revenue if and only if $J(\theta)$ is increasing.
If $J(\theta)$ is not monotonic, define $\tilde{J}(\theta)=\int_{0}^{\theta}J(x)\mathop{}\!\mathrm{d}F(x)$ and let $\operatorname{conv}\tilde{J}$ denote the convex hull of $\tilde{J}$. Then, the revenue-maximizing mechanism pools types into the same positional good level if $\operatorname{conv}\tilde{J}$ is affine and separates types if $\operatorname{conv}\tilde{J}=\tilde{J}$.
@@
GW’s Proposition 8 provides a sufficient condition for the seller’s revenue to be strictly increasing in the number of priority classes: $F$ satisfies the IFR property.
@@
Effort maximization in MSS’s model is the same as revenue maximization because they assume linear effort costs (and no exclusion). In Theorem 4, they also provide a sufficient condition for full separation : $F$ satisfies the IFR property.
@@
Proof sketch.
Because $J^{\prime}(\bar{\theta})=2>0$, $J(\theta)$ is increasing in a neighborhood of $\bar{\theta}$, so by Proposition 2 in KMS, types in this neighborhood are separated. ∎
@@
The seller can obtain at least half the maximum revenue by selling a single positional good in addition to a free low-level position.
@@
By offering one positional good in addition to a free low-level position, the seller’s maximum revenue is
@@
Figure 4 plots the revenue curve $R(\tau)=(1-\tau)F^{-1}(\tau)$ in the quantile space $\tau=F(\theta)$. Assume $v(\theta)=0$ for simplicity. Similar to Section 3.1.1, (i) the area under the revenue curve (blue area in the left panel) represents the revenue from the revenue-maximizing mechanism—i.e., $s^{*}(\theta)=F(\theta)$, (ii) the red triangle in the right panel represents the revenue from having two levels of positional goods—a higher level offered at price $F(\theta^{*}_{0})$ in addition to a lower level offered for free, and (iii) the entire box of either panel (which has an area of $R(\tau^{*}_{0})$) represents the maximum revenue from selling nonpositional goods in the auxiliary problem. It is straightforward that the red area (ii) is larger than half of the entire box and thus strictly larger than half of the blue area (i).
@@
If $F(\theta)=1-\exp(-\lambda\theta)$ where $\lambda>0$, offering a single good above the free lowest level can obtain $73.6\%$ of the maximum revenue.
@@
If $F(\theta)=\theta$ on $[0,1]$, offering a single good above the free lowest level can obtain $75\%$ of the maximum revenue.
@@
If $F(\theta)=\theta^{\beta}$ on $[0,1]$ where $\beta>0$, offering a single good above the free lowest level can obtain
@@ A.2 Consumer Surplus
I focus on the case where the price (or effort) is nonnegative, which pin down p(0)=v(0)U(0)=0p(0)=v(0)-U(0)=0.
@@
If both $J(\theta)=\theta-\frac{1-F(\theta)}{f(\theta)}$ and $\frac{1-F(\theta)}{f(\theta)}$ are increasing (e.g., exponential and Pareto distributions), a finer partition (in terms of majorization) can increase both the seller’s revenue and consumer surplus. Thus, full separation maximizes both the revenue and consumer surplus.
@@
[…]
@@
Define $I(\theta)=\int_{0}^{\theta}\frac{1-F(x)}{f(x)}\mathop{}\!\mathrm{d}F(x)$. Then, the condition in the proposition is equivalent to $I(\theta)\geq I(\bar{\theta})F(\theta)=\operatorname{\mathbf{E}}[\theta]F(\theta)$ (i.e., $I(\theta)$ lies above the line connecting $I(0)=0$ and $I(\bar{\theta})=\operatorname{\mathbf{E}}[\theta]$ in the quantile space). Therefore, $I(\bar{\theta})F(\theta)=\operatorname{conv}I$ (the largest convex function that lies below $I$). By Proposition 2 in KMS, this condition is necessary and sufficient for total pooling to be welfare-maximizing. ∎
@@
$\max_{s\in\operatorname{\text{MPS}}(F)}\int_{0}^{\theta}\frac{s(x)}{f(x)}\mathop{}\!\mathrm{d}F(x)+v(\theta)$ (A.9)
The following proposition provides a necessary and sufficient condition for total pooling to maximize $U(\theta)$ for all $\theta\in[0,\bar{\theta}]$.
@@
Note that $\int_{0}^{\theta}1/f(x)\mathop{}\!\mathrm{d}F(x)=\theta$, and the condition $F(\theta)\leq\theta/\bar{\theta}$ is equivalent to $\int_{0}^{\theta}(1/f(x)-\bar{\theta})\mathop{}\!\mathrm{d}F(x)\geq 0$ for all $\theta\in[0,\bar{\theta}]$. Similar to the proof of Proposition A.5, one can show that this condition is necessary and sufficient for $s(\theta)=1/2$ to maximize $U(\theta)$ for all $\theta\in[0,\bar{\theta}]$. ∎
@@ Appendix B Proofs
B.1 Proof of Theorem 1
Proof.
[…]
and (iii) fully separating otherwise, where $s(\theta)=F(\theta)$.
Hence, by Proposition 1 (Choquet’s Theorem) in KMS, there exists a probability measure $\lambda_{s}$ supported on the extreme points of $\operatorname{\text{MPS}}_{w}(F)$ such that $s=\operatorname{\mathbf{E}}[\tilde{s}\mid\tilde{s}\sim\lambda_{s}]$. Thus, $s$ can be induced by randomizing over allocations that each induces extreme points of $\operatorname{\text{MPS}}_{w}(F)$. ∎
B.2 Proofs of Section 3
B.2.1 Proofs of Section 3.1
@@ B.2 Proofs of Section 3
The following lemma follows from applying Theorem 4 (Fan-Lorentz) in KMS to
@@
If $J(\theta)$ is not monotonic, the revenue-maximizing mechanism pools types into the same positional good level if $\operatorname{conv}\tilde{J}$ is affine and separates types if $\operatorname{conv}\tilde{J}=\tilde{J}$. ∎
@@
Define $J_{v}(\theta)=v(\theta)-\frac{1-F(\theta)}{f(\theta)}v^{\prime}(\theta)$. Under IFR, $J_{v}(\theta)$ is increasing because $v(\theta),v^{\prime}(\theta)\geq 0$ and $v^{\prime\prime}(\theta)\leq 0$. Define
@@
Define $\theta_{1}=J^{-1}(0)$. We have
@@
Denote by $p$ the price of the positional good. The cutoff type $\theta(p)$ indifferent between buying and not buying is given by
@@
Because $v^{\prime}(\theta)\geq 0$, all types above $\theta(p)$ will buy the good. Thus, by selling a single good, the seller’s maximum revenue is
@@
Consider the auxiliary screening problem of selling an indivisible item of value to one buyer, in which a standard result implies a posted-price mechanism is optimal (see Börgers, 2015, Proposition 2.5). Formally, denote by $q$ the probability the item is allocated to $\theta$, and let $\mathcal{M}=\{q\colon[0,\bar{\theta}]\to[0,1]\mid q\text{ increasing}\}$. The agent’s payoff is $\theta q-p$. Then, the standard result implies that $\max_{q\in\mathcal{M}}\int_{0}^{\bar{\theta}}J(\theta)q(\theta)\mathop{}\!\mathrm{d}F(\theta)=\max_{p}p(1-F(p))$, and any maximizer $q^{*}$ must be an extreme point of $\mathcal{M}$ (i.e., a posted-price mechanism).
@@
(i) The following lemma follows from applying Theorem 4 (Fan-Lorentz) in KMS to
$W(\theta_{0})=\int_{\theta_{0}}^{\bar{\theta}}\left(\frac{1-F(\theta)}{f(\theta)}\right)(s(\theta)+v^{\prime}(\theta))+U(\theta_{0})\mathop{}\!\mathrm{d}F(\theta)$ (B.5)
@@
If $s(\theta)=s_{0}$ is constant on $\theta_{0},\theta_{1}]$ for some $\theta_{1}\in(\theta_{0},\bar{\theta}]$ (and $s(\theta)>s_{0}$ for $\theta>\theta_{1}$), then the consumer surplus is
@@
$\begin{aligned} W^{\prime}(\theta_{0})=-({1-F(\theta_{0})})(v^{\prime}(\theta_{0})+F(\theta_{1})/2+F(\theta_{0})/2)+\frac{f(\theta_{0})}{2}\int_{\theta_{0}}^{\theta_{1}}({1-F(\theta)})\mathop{}\!\mathrm{d}\theta.\end{aligned}$
Define
$z(\theta_{0})\equiv-\frac{W^{\prime}(\theta_{0})}{1-F(\theta_{0})}=v^{\prime}(\theta_{0})+F(\theta_{1})/2+F(\theta_{0})/2-\frac{f(\theta_{0})}{1-F(\theta_{0})}\int_{\theta_{0}}^{\theta_{1}}\frac{1-F(\theta)}{2}\mathop{}\!\mathrm{d}\theta$
and $\psi(\theta)=z(\theta)-F(\theta)-v^{\prime}(\theta)$. Note that $\psi(\theta_{1})=0$, and under IFR,
$\psi^{\prime}(\theta)=-\left[\frac{f(\theta)}{1-F(\theta)}\right]^{\prime}\int_{\theta}^{\theta_{1}}\frac{1-F(\theta)}{2}\mathop{}\!\mathrm{d}\theta\leq 0.$
Therefore, $\phi(\theta_{0})\geq 0$ for all $\theta_{0}\leq\theta_{1}$. It follows that $z(\theta_{0})\geq F(\theta_{0})+v^{\prime}(\theta_{0})\geq 0$ and thus $W^{\prime}(\theta_{0})=-z(\theta_{0})(1-F(\theta_{0}))\leq 0$. Hence, decreasing $\theta_{0}$ increases the consumer surplus $W(\theta_{0})$. ∎
@@
$\begin{aligned} W^{\prime}(\theta_{0})=-({1-F(\theta_{0})})(v^{\prime}(\theta_{0})+1/2+F(\theta_{0})/2)+\frac{f(\theta_{0})}{2}\int_{\theta_{0}}^{\bar{\theta}}({1-F(\theta)})\mathop{}\!\mathrm{d}\theta\end{aligned}$ (B.8)
The derivative is the same as in the proof of Proposition 4, so the same argument implies that $W^{\prime}(\theta_{0})\leq 0$ for all $\theta_{0}>0$ and hence no exclusion is optimal (i.e., $\theta^{*}_{0}=0$).